the area of a triangle is 27 square inches. if the height is 3 inches less than two times the base, determine the height and base of the triangle
Answer by user_dude2008(1862)

A=(bh)/2
27=(b(2b-3))/2
54 = 2b^2-3b

2b^2-3b-54=0

(2b+9)(b-6)=0

2b+9=0 or b-6=0
b=-9/2 or b=6

Answer: base = 6, height = 9